OpenStudy (bangbang):

For the values a = 2.9 which is a leg of a right triangle, and value c = 4.2, which is the hypotenuse, find the length of the other leg, b, to the nearest tenth

OpenStudy (johnweldon1993):

Right triangle...you can use the Pythagorean Theorem a² + b² = c² 'c' is your hypotenuse ....'a' and 'b' are the OTHER 2 legs....we have the hypotenuse...and one of the other legs....we need to solve for 'b' \[b^2 = c^2 - a^2\] \[b = \sqrt{c^2 - a^2}\] input your values of 'a' and 'c' and find 'b'